OpenStudy (anonymous):

re edit question no.2

OpenStudy (anonymous):

find the derivative of y = x^2 cos x- 2xsinx - 2cos x with respect to x. a. x^2 sinx b.-x^2 cosx c.x^2 cosx d.-x^2 sinx

OpenStudy (anonymous):

i got the answer here letter D

OpenStudy (nenadmatematika):

=2xcosx-x^2sinx-(2sinx+2xcosx)+2sinx=2xcosx-x^2sinx-2sinx-2xcosx+2sinx=-x^2sinx so D

OpenStudy (anonymous):

let's define y / x = alpha then sin (alpha) = 1/alpha this equation determines alpha as a constant, therefore the relationship between y and x is linear and we have y = alpha x and subsequently y' = alpha however we defined alpha to be y/x therefore y' = y/x <<<answer from re edit question 1